OpenStudy (ttp):

The point Negative five, two would not fall on the line for which of the following representations? Negative three y equals x minus one. Six y minus two equals negative two x. five plus y equals two x minus four. Eleven minus six x equals thirteen y plus fifteen.

OpenStudy (ttp):

@e.cociuba

OpenStudy (ttp):

@Easyaspi314

OpenStudy (anonymous):

plug in (x, y) = (-5, 2) in every equation. find the one for which it doesn't hold true. first one: -3y = x - 1 plug in (-5, 2) which ix x = -5 and y =2 -3(2) = (-5) - 1 this equation holds true [ -6 = -6 ]. therefore the point (-5, 2) IS on that line

OpenStudy (ttp):

So i'm confused what one

OpenStudy (anonymous):

you just want the answer?

OpenStudy (ttp):

Can you help me get the answer?

OpenStudy (anonymous):

B: 6y - 2 = -2x plug in (-5, 2) 6(2) - 2 = -2(-5) both sides equal 10 so this is true. this is not the answer (read question) C: 5 + y = 2x - 4 5 + (2) = 2(-5) - 4 7 = -14 This isn't true. so the point (-5, 2) is NOT on the line 5+ y = 2x - 4 [y = 2x - 9] answer is C..